Constants Part A A point charge q1-15.00 μC is held fixed in space. From a horizontal distance of 6.00 cm, a small sphere with mass 4.00 x 10-3 kg and charge What is the acceleration of the sphere at the instant when its speed is 26.0 m/s? Express your answer with the appropriate units. +2.00 μC is fired toward the fixed charge with an initial speed of 41.0 m/s Gravity can be neglected a: 1 Value Units
Answer #1

Solution- then distance is travels is y

0.5 mV^2 = Kq1q2*(1/r1-1/r)

1/r1-1/r = 0.5*4*10^-3*41^2/9*10^9*15*10^-6*2*10^-6

=12.45 m

1/r=(1/0.06)+12.45

r=3.43 cm

So,

v^2-u^2=2as

So, a= 41^2 -26^2 / (2*0.0343)

a=14.650*10^3 m/s^2

I hope help you !!